OpenStudy (anonymous):

950 cm^3=0.95_______

OpenStudy (anonymous):

dm^3

OpenStudy (anonymous):

how did you get that answer?

OpenStudy (anonymous):

950 / 1,000 = 0.95 A unit type between meter and kilometer.

OpenStudy (anonymous):

(1cm)^3 = (0.1dm)^3 1 cm^3 = 0.001dm^3 950 cm^3 = 0.001(950)

OpenStudy (anonymous):

He's right its Decimeter.

OpenStudy (anonymous):

950cm^3 = .95dm or 950cm^3 = .95dm^3 Either one of those

OpenStudy (anonymous):

Thank you I'm not really good at math so I need all the help I can get

OpenStudy (anonymous):

No prob.

OpenStudy (anonymous):

@tkins89 what grade are you in

OpenStudy (anonymous):

@ surdawi im in college

OpenStudy (anonymous):

since you are in college that should be easy stuff anyways do you understand what i did

OpenStudy (anonymous):

Im not a math person!!! and yes I did thank you very much

OpenStudy (anonymous):

im in college too, so if you need help i might be able to help you